### v3

# Meta-Generalized-Gradient Approximation made Magnetic

Abstract
The Jacob’s ladder of density functional theory (DFT) proposes the compelling view that by extending the form of successful approximations — being guided by exact conditions and selected (least empirical) norms — upper rungs will do better than the lower, thus allowing to balance accuracy and computational effort. Meta-generalized-gradient-approximations (MGGAs) belong to the last rung of the semi-local approximations before hybridization with non-local wave function theories. Among the MGGAs, the Strongly Constrained and Appropriately Normed Approximation (SCAN) greatly improves upon GGAs from the lower rung. But the over magnetized solutions of SCAN make GGAs more reliable for magnetism. Here, we provide a solution that satisfies the most pressing desiderata for a density functional approximations for ferromagnetic, antiferromagnetic and non-collinear states. The approach is available in an implementation in the Crystal electronic structure package.
pacs:
71.15.Mb, 71.15Rf, 31.15.E-
Introduction.  Despite nearly one hundred years of quantum physics, magnetism continues to challenge our understanding of electron correlations. An unsettling example is given by the recently developed strongly-constrained and appropriately normed (SCAN) meta-generalized gradient approximation (MGGA) for density functional theory (DFT).Sun et al. (2015, 2016)
SCAN satisfies a large set of exact constraints, from known properties of the exact exchange-correlation (xc) functional and gives accurate results at a computational cost of about a generalized-gradient-approximation (GGA). Thus, it can be considered a modern milestone from which it is possible to develop more accurate approximations.Aschebrock and Kümmel (2019); Lebeda et al. (2024)
But SCAN does not reproduce correctly the stability and properties of prototypical ferromagnets (FM) like Fe. Fu and Singh (2018); Ekholm et al. (2018); Fu and Singh (2019); Mejía-Rodríguez and Trickey (2019); Tran et al. (2020) Overestimated tendency toward magnetism and exaggeration of magnetic energies are obtained for transition metal systems. Extension of the analysis to antiferromagnetic (AFM) solids has shown that modifications of MGGAs yielding an improved description of FM metals produce a deterioration in AFM insulators, and vice versa.Tran et al. (2020) The status is such that MGGAs yield no advantage over lower rungs of the semi-local approximation.Romero and Verstraete (2018); Tran et al. (2020)
Here, we argue that the problem has persisted because the most important exact constraint for dealing with magnetism has not guided most of the DFT developments thus far. We resolve this issue by switching from collinear spin-DFTvon Barth and Hedin (1972); Gunnarsson and Lundqvist (1976) (the current standard methodology) to the fully non-collinear (nc) Vignale-Rasolt-Bencheikh spin-current DFT (SCDFT) – a complete DFT theory for dealing with magnetism within the framework of the two-component Pauli equation.Vignale and Rasolt (1988); Bencheikh (2003) As will be shown, this extended framework can also solve problems for collinear DFT.
In SCDFT, the most distinctive property of the exact xc energy functional is the form invariance under U(1)$\times$×SU(2) gauge transformations.Vignale and Rasolt (1988); Bencheikh (2003); Pittalis et al. (2017) These transformations rotate spinors as well as their phase at each point in space [see Eq. (11) below]. Failing to satisfy this invariance leads to functional forms that do not necessarily distinguish between physical variations of the spin density and those that arise from a mere change of reference frame. The gauge invariance of the exact xc-energy functional is thus the most important exact constraint for guiding the construction of magnetic approximations.
Working within the general SCDFT, extra building blocks for the construction of extended approximate functional forms can be gained.Pittalis et al. (2017) In this work, we shall see, the collinear isoorbital indicators used in the SCAN (more below), which treat spin-up and spin-down channels separately, yield unbalanced magnetic solutions, Mejía-Rodríguez and Trickey (2019); Tran et al. (2020) and can be replaced by a unified non-collinear quantity normed by the gauge principle. As a result, the major issues observed in the collinear SCAN solutions get resolved, and the revised functional form is seamlessly extended to the more general non-collinear regime.
From DFT to collinear Spin-DFT.  DFT can in principle find the particle density and total energy of real interacting ground states by solving (self-consistently) an auxiliary non-interacting problem
$\left\{-\frac{1}{2}\nabla^{2}+v_{\rm ext}+\int d^{3}r^{\prime}\frac{n({\bf r}^{\prime})}{|{\bf r}-{\bf r}^{\prime}|}+v_{{\rm xc}}\right\}\varphi_{i}=\varepsilon_{i}\varphi_{i}\;,$ (1)
which is the celebrated Kohn-Sham equation. This is a Schrödinger-like equation that describes single-particle orbitals corresponding to the actual external potential (the second term on the l.h.s.) added to the Hartree potential (the third term on the l.h.s.) and the xc potential, $v_{\rm xc}$. The latter potential is such that the actual particle density is retrieved $n=\sum^{occ}_{i}|\varphi_{i}|^{2}$not (a) Correspondingly, the energy of the ground state of the real system $E$ is thus obtained by
$E=T_{s}+\int d^{3}r~{}n~{}v_{\rm ext}+E_{\rm H}[n]+E_{\rm xc}[n]\;,$ (2)
where $T_{s}=\frac{1}{2}\sum_{i}^{\rm occ}\int d^{3}r~{}|\nabla\varphi_{i}|^{2}$ is the Kohn-Sham kinetic energy, $E_{\rm H}[n]=\int\int d^{3}rd^{3}r^{\prime}\frac{n({\bf r})n({\bf r}^{\prime})}{2|{\bf r}-{\bf r}^{\prime}|}$ is the Hartree energy and $E_{\rm xc}[n]$ is the xc energy functional. For retrieving the exact total energy of the actual state, $E_{\rm xc}[n]$ accounts for the electron-electron energy beyond the classical evaluation obtained via $E_{\rm H}[n]$ as well as the kinetic energies beyond $T_{s}$. Also notice that $v_{\rm xc}=\delta E_{\rm xc}/\delta n$. All that is needed to make the above approach practical is a density functional approximation (DFA) for $E_{\rm xc}[n]$. Solutions to $n$ and $E$ can thus be found in an iterative fashion.
The first and simplest $E^{\rm DFA}_{\rm xc}[n]$ is the local density approximation (LDA) Kohn and Sham (1965)
$E_{\rm xc}^{\rm LDA}[n]=\int d^{3}r~{}n~{}\epsilon^{\rm unif}_{\rm xc}(n)\;,$ (3)
where the xc energy density is approximated locally with that of the uniform gas, $\epsilon^{\rm unif}_{\rm xc}(n)$, for which analytical and accurate numerical reference data are available. To capture the variations of more realistic states, a GGA uses information retrieved from the gradients of the density via $s=|\nabla n|/\left[2(3\pi^{2})^{1/3}n^{4/3}\right]$
$E_{\rm xc}^{\rm GGA}[n]=\int d^{3}r~{}n~{}\epsilon^{\rm unif}_{\rm xc}(n)~{}F_{\rm xc}^{\rm GGA}(n,s)\;,$ (4)
where $F_{\rm xc}^{\rm GGA}(n,s)$ is the xc enhancement factor that connects the LDA energy densities to the GGA energy densities. For slowly varying density, the LDA is retrieved (i.e. $F_{\rm xc}^{\rm GGA}(n,0)=1$). The most popular choice for the enhancement factor is the one by Perdew, Burke and Ernzerhof (PBE) $F_{\rm xc}^{\rm PBE}$. Perdew et al. (1996)
MGGAs Perdew et al. (1999); Tao et al. (2003); Sun et al. (2015) introduce an extra dependence beyond the electron density by employing
$E_{\rm xc}^{\rm MGGA}[n]=\int d^{3}r~{}n~{}\epsilon^{\rm unif}_{\rm xc}(n)~{}F^{\rm MGGA}_{\rm xc}(n,s,\tau)\;,$ (5)
where $\tau=1/2\sum_{i}^{\mathrm{occ}}|\nabla\varphi_{i}|^{2}$ is the (KS) kinetic energy density. Kaplan and Perdew (2022) Here we focus on those MGGAs that depend on $\tau$ via
$\alpha:=\frac{\tau-\frac{|\nabla n|^{2}}{8n}}{\tau_{\rm unif}}\;,$ (6)
where $\tau_{\rm unif}=\frac{3}{10}(3\pi^{2})^{2/3}\,n^{5/3}$ is the kinetic energy density of the uniform gas at density $n$. In particular, the SCAN form is defined by setting $F_{\mathrm{xc}}^{\mathrm{SCAN}}(n,s,\alpha)=h_{\mathrm{xc}}^{1}(n,s)+\left[h_{\mathrm{xc}}^{0}(n,s)-h_{\mathrm{xc}}^{1}(n,s)\right]f_{\mathrm{xc}}(\alpha)$ where $f_{\mathrm{xc}}(\alpha)$ interpolates between the semi-local xc enhancement factors for single-orbital densities ($h_{\mathrm{xc}}^{0}$) and for slowly-varying densities ($h_{\mathrm{xc}}^{1}$). This modeling exploits the fact that $\alpha$ [see Eq. (6) ] tends to zero in the iso-orbital limit (i.e., when the density is dominated by a single occupied orbital), tends to one in the uniform-density limit, and becomes $\gg 1$ in regions of low density between closed shells
For solving ground states involving collinear spin polarization either induced by external Zeeman couplings or spontaneously via magnetic phase transitions, the standard approach is to switch from DFT to Spin-DFT.von Barth and Hedin (1972); Gunnarsson and Lundqvist (1976) The method requires to solve a spin-dependent, single-orbital problem
$\left\{-\frac{1}{2}\nabla^{2}+v_{{\rm ext},\sigma}+v_{{\rm Hxc},\sigma}\right\}\varphi_{i,\sigma}=\varepsilon_{i,\sigma}\varphi_{i,\sigma}\;.$ (7)
Here, $E_{\rm H}[n]$ and corresponding $v_{\rm H}$ are unchanged but the spin-dependent xc-potential, $v_{{\rm xc},\sigma}=\delta E_{\rm xc}/\delta n_{\sigma}$, is generated from the functional $E_{\rm xc}[n_{\uparrow},n_{\downarrow}]$, which depends both on $n_{\uparrow}$ and $n_{\uparrow}$ and not just on their sum $n=\sum_{\sigma}n_{\sigma}$.
Extended forms for exchange can be generated from given DFT approximations using the collinear spin decomposition for exchange not (b)
$E_{\rm x}^{\rm SDFA}[n_{\uparrow},n_{\downarrow}]=\frac{1}{2}E_{\rm x}^{\rm DFA}[2n_{\uparrow}]+\frac{1}{2}E_{\rm x}^{\rm DFA}[2n_{\downarrow}]\;.$ (8)
For correlation, we must use
$E_{{\rm c}}^{\rm SDFA}[n_{\uparrow},n_{\downarrow}]=\int d^{3}r~{}n~{}\varepsilon^{\rm unif}_{\rm c}(n,\zeta)~{}F^{\rm SDFA}_{\rm c}(n,\zeta,s,\alpha)\;.$ (9)
Unlike exchange, the correlation functional depends in a non separable fashion on $n_{\uparrow}$ and $n_{\downarrow}$ through the spin polarization (density) $\zeta=|(n_{\uparrow}-n_{\downarrow})|/(n_{\uparrow}+n_{\downarrow})$; $\alpha=(\tau-\frac{|\nabla n|^{2}}{8n})/\tau_{\rm unif}d_{s}$ is the spin-summed iso-orbital indicator with $\tau=1/2\sum_{i,\sigma}^{\mathrm{occ}}|\nabla\varphi_{i}|^{2}$ and $d_{s}=\frac{1}{2}[(1+\zeta)^{5/3}+(1-\zeta)^{5/3}]$. Also note that the exchange functional of Eq. (8), $E_{\rm x}^{\rm DFA}[n_{\uparrow},n_{\downarrow}]$, can be equivalently reexpressed as $E_{\rm x}^{\rm DFA}[n,\zeta]$ in terms of the total density $n$ and polarization $\zeta$.
Issues.  Notice that $\alpha$ does not carry any explicit information on $\nabla\zeta$. Therefore, the above expressions for GGAs and MGGAs come with a striking asymmetry. While the gradients of $\zeta$ are included in the exchange functional indirectly via the spin-scaling relation [i.e., Eq. (8)], the correlation form has no dependence on $\nabla\zeta$.
This asymmetry reflects the lack of invariance of the approximate forms under U(1)$\times$×SU(2) gauge transformations. Although this deficiency might not be perceived as a problem as long as we consider scenarios that are not gauge symmetric — such as FMs or AFMs — one should not confuse the symmetry properties of particular solutions with those of the exact functional. To sort the latter out, we must enlarge our modeling framework.
Switching to SCDFT.  Let us account not only for a scalar multiplicative potential $v_{\rm ext}$ and a multiplicative magnetic field ${B}^{a}$, but also for a (particle-) vector potential ${\bf A}$ and a spin-vector potential ${\bf A}^{a}$not (c) While ${\bf A}$ is useful to represent the vector potential of an external magnetic field, ${\bf A}^{a}$ is useful to represent the (one-body) spin-orbit couplings in the system. Vignale and Rasolt (1988); Bencheikh (2003)
The KS equations in SCDFT have the form of single-particle Pauli equations Vignale and Rasolt (1988); Bencheikh (2003)
$\left[\frac{1}{2}\left(-i\nabla+\frac{1}{c}{\mathbfcal{A}}_{\rm{\tiny\tiny KS}}\right)^{2}+{\cal V}_{\rm{\tiny\tiny KS}}\right]\Phi_{k}=\varepsilon_{k}\Phi_{k}\;,$ (10)
where $\Phi_{k}$ are two-component spinors, ${\mathbfcal{A}}_{\rm{\tiny\tiny KS}}=\left({\bf A}+{\bf A}_{{\rm xc}}\right)+\sigma^{a}\left({\bf A}^{a}+{\bf A}^{a}_{{\rm xc}}\right)$, ${\cal V}_{\rm{\tiny\tiny KS}}=\left(v+v_{\rm H}+v_{\rm xc}\right)+\sigma^{a}\left(B^{a}+B^{a}_{\rm xc}\right)+\frac{1}{2c^{2}}\left[\left({\bf A}+\sigma^{a}{\bf A}^{a}\right)^{2}-{\mathbfcal A}^{2}_{\rm{\tiny\tiny KS}}\right]$, in which $\sigma^{a}$ denotes the component of the vector of Pauli matrices, ${\frac{1}{c}}{\bf A}_{\rm xc}=\delta E_{\rm xc}/\delta{\bf j}$ is an Abelian xc-vector potential, ${\frac{1}{c}}{\bf A}^{a}_{\rm xc}=\delta E_{\rm xc}/\delta{\bf J}^{a}$ is the $a$-th component of a non-Abelian xc-vector potential, ${B}^{a}_{\rm xc}=\delta E_{\rm xc}/\delta{m^{a}}$ is the $a$-th component of a xc-magnetic-like potential, $v_{\rm xc}=\delta E_{\rm xc}/\delta n$ is a xc-scalar potential, and $v_{\rm H}$ is the usual Hartree potential. The KS densities are obtained from the (occupied) two-component KS spinors as follows: $n=\sum_{k}^{\rm occ}\Phi^{\dagger}_{k}\Phi_{k}$, ${\vec{m}}=\sum_{k}^{\rm occ}\Phi^{\dagger}_{k}\;{\vec{\sigma}}\;\Phi_{k}$, ${\bf j}=\frac{1}{2i}\sum_{k}^{\rm occ}\Phi^{\dagger}_{k}\left[\nabla\Phi_{k}\right]-\left[\nabla\Phi^{\dagger}_{k}\right]\Phi_{k}$, and ${\vec{\bf J}}=\frac{1}{2i}\sum_{k}^{\rm occ}\Phi^{\dagger}_{k}{\vec{\sigma}}\left[\nabla\Phi_{k}\right]-\left[\nabla\Phi^{\dagger}_{k}\right]{\vec{\sigma}}\Phi_{k}$.
Because $E_{\rm xc}[n,{\vec{m}},{\bf j},{\vec{\bf J}}]$ must yield the exact interacting densities via the KS densities, it must also be form invariant under (local) U(1)$\times$×SU(2) gauge transformations.Vignale and Rasolt (1988); Bencheikh (2003); Pittalis et al. (2017) In detail, such a transformation acts on spinors as follows
${\Phi}^{\prime}({\bf r})=\exp\left[\frac{i}{c}\chi({\bf r}){\mathrm{I}}+\frac{i}{c}{\lambda^{a}}({\bf r}){\sigma^{a}}\right]{\Phi}({\bf r})\;,$ (11)
where I is the $2\times 2$ identity and $\chi$ and $\lambda^{a}$ are arbitrary real-valued functions.
We recall from Eq. (5) that MGGAs, such as SCAN, explicitly depend on the KS kinetic energy density $\tau$. For dealing with nc solutions, the kinetic energy density $\tau$ must be expressed in terms of the occupied two-component KS spinors $\tau_{\rm nc}=1/2\sum_{k}^{\mathrm{occ}}\Big{(}\partial_{\mu}\Phi^{\dagger}_{k}\Big{)}\Big{(}\partial_{\mu}\Phi_{k}\Big{)}$. But, it can be directly verified that $\tau_{\rm nc}$ is not gauge invariant.Pittalis et al. (2017) Thus, $\alpha_{\rm nc}$ obtained by replacing $\tau$ with $\tau_{\rm nc}$ in Eq. (6) is not invariant either. The corresponding extension of the SCAN functional by letting $\alpha\to\alpha_{\rm nc}$ would therefore lack the most fundamental property for dealing with magnetism. Fortunately, this problem can be fixed by the replacement ${\tau}_{\rm nc}\rightarrow\widetilde{\tau}_{\rm nc}$ withPittalis et al. (2017); Desmarais et al. (2024a)
$\widetilde{\tau}_{\rm nc}=\tau_{\rm nc}+\frac{{m}^{a}{\tau}_{\rm nc}^{a}}{n}+\frac{\nabla{m}^{a}\cdot\nabla{m}^{a}}{8n}-\frac{{\bf j}\cdot{\bf j}}{2n}-\frac{{\bf J}^{a}\cdot{\bf J}^{a}}{2n}\;,$ (12)
where ${\tau}^{a}_{\rm nc}=1/2\sum_{k}^{\rm occ}\Big{(}\partial_{\mu}\Phi^{\dagger}_{k}\Big{)}{\sigma}^{a}\Big{(}\partial_{\mu}\Phi_{k}\Big{)}\;$ is the spin-kinetic energy density.
Eq. (12) not only includes $\nabla\vec{m}$ and the (spin-)paramagnetic currents but, for consistency with the exact gauge-invariance requirements, also the coupling between ${\vec{m}}$ and ${\vec{\tau}}_{\rm nc}$.
Eq. (12) includes the well-known particle current corrections that ensure the U(1) invariance of the MGGA, which has been advanced both in practice and in theory.Dobson (1993); Becke (2002); Tao and Perdew (2005); Pittalis et al. (2006, 2007); Räsänen et al. (2009); Furness et al. (2015); Tellgren et al. (2018); Laestadius et al. (2019); Pemberton et al. (2022) It also includes the spin current corrections that were recently shown to ensure the SU(2) invariance of the MGGA for time-reversal symmetric spin-orbit coupling — the JSCAN.Desmarais et al. (2024b) Eq. (12) was also at the center of the derivation of a non-collinear ELF.Desmarais et al. (2024a) But it did not find exploitation in energy functional approximations to determine magnetic solutions.
We recall that the problems of overmagnetization by SCAN were imputed, in Refs. Mejía-Rodríguez and Trickey, 2019; Tran et al., 2020, to the different behaviour of the isoorbital indicators in the exchange of SCAN for the two spin channels. Below, we solve these problems by invoking a single $\widetilde{\alpha}$ for both spin channels — normed by the gauge principle. For capturing energies rather than just bond pictures, the following gauge invariant extension of SCAN suggests itself
$E^{\rm nc}_{x}[n,\vec{m},\mathbf{j},\vec{\mathbf{J}}]=\int d^{3}r~{}n~{}\varepsilon^{\rm unif}_{x}\left(n,\zeta_{\rm nc}\right)~{}F^{\rm SCAN}_{x}\left(n,s,\widetilde{\alpha}_{\rm nc}\right)\;,$ (13a)
and
$E^{\rm nc}_{c}[n,\vec{m},\mathbf{j},\vec{\mathbf{J}}]=\int d^{3}r~{}n~{}\varepsilon^{\rm unif}_{\rm c}(n,\zeta_{\rm nc})~{}F^{\rm SCAN}_{\rm c}(n,\zeta_{\rm nc},s,\widetilde{\alpha}_{\rm nc})\;,$ (13b)
where $\zeta_{\rm nc}\equiv|\vec{m}|/n$, and
$\widetilde{\alpha}_{\rm nc}\equiv\frac{2n\widetilde{\tau}_{\rm nc}-\frac{(\nabla n)\cdot(\nabla n)}{4}}{\left(n+|\vec{m}|\right)\tau_{\rm unif}^{+}+\left(n-|\vec{m}|\right)\tau_{\rm unif}^{-}}$ (14)
in which $\tau_{\rm unif}^{\pm}\equiv 3/10(3\pi^{2})^{2/3}(n\pm|\vec{m}|)^{5/3}$, and all uniform gas quantities have been upgraded to an nc form by locally quantizing spin along the axis defined by the spin density $\vec{m}$.Kübler et al. (1988); not (d) The derivation of Eqs. (14) and (15) is provided in the Supplementary Material (this includes Ref.s Giuliani and Vignale, 2005; Oliver and Perdew, 1979; Lehtola et al., 2020; Desmarais et al., 2021; Vilela Oliveira et al., 2019; Peterson et al., 2007; Dolg, 2005; Balabanov and Peterson, 2005; Wilson et al., 1996; Campetella et al., 2020; Saunders et al., 1992; Towler et al., 1996; Doll et al., 2001, 2006; Desmarais et al., 2023; Dachsel et al., 1999; Yang et al., 2018; Wenzel et al., 2022; El Idrissi et al., 1991; Jones et al., 2022; Huber, 1979).ESI ; not (e)
For self-consistent collinear solutions, a modified SCAN functional (named mSCAN) is straightforwardly obtained solely in terms of $n^{\sigma}$, $\varepsilon^{\rm unif}_{\rm x/c}(n^{\uparrow},n^{\downarrow})$, and
$\widetilde{\alpha}\equiv\frac{n^{\uparrow}\tau^{\uparrow}+n^{\downarrow}\tau^{\downarrow}-\frac{\left(\nabla n^{\uparrow}\right)\cdot\left(\nabla n^{\downarrow}\right)}{4}}{n^{\uparrow}\tau^{\uparrow}_{\rm unif}+n^{\downarrow}\tau^{\downarrow}_{\rm unif}}\;,$ (15)
where $\tau^{\sigma}_{\rm unif}=3/10(6\pi^{2})^{2/3}(n^{\sigma})^{5/3}$.not (f) Finally, the mSCAN reduces to the SCAN on closed-shell solutions (more below).
Notable magnetic SCAN failures — resolved. Let’s scrutinize the novel collinear solutions. We implemented the mSCAN functional in a developer’s version of the Crystal23 program.Erba et al. (2023) We study current-less, collinear magnets. Thus, the single-particle equation of SCDFT given in Eq. (10) is considerably simplified, as the vector potentials ${\bf A}_{\rm xc}$ and ${\bf A}^{a}_{\rm xc}$ vanish, and we must self-consistently solve the simpler Eq. (7) in a basis of Bloch orbitals $\varphi_{i,\sigma,\mathbf{k}}$ (with wavevector $\mathbf{k}$) formed from a linear combination of atom-centered spherical Gaussian-type functions (GTF).Desmarais et al. (2018) In Eq. (7) for mSCAN, in a generalized KS context,Seidl et al. (1996); Desmarais et al. (2024c) the spin-dependent xc potential $v_{{\rm xc},\sigma}$ readsNeumann et al. (1996)
$v_{{\rm xc},\sigma}=\left.\frac{\delta E^{\rm mSCAN}_{xc}}{\delta n^{\sigma}}\right|_{\tau^{\sigma}}-\nabla\cdot\frac{1}{2}\left.\frac{\delta E^{\rm mSCAN}_{xc}}{\delta\tau^{\sigma}}\right|_{n^{\sigma}}\cdot\nabla$ (16)
We solve Eqs. (7) and (16). Computational details are provided in the SM.
Before applying mSCAN to prototypical magnets, notice that the x-mSCAN does not fulfill the (strictly collinear) spin decomposition relation of Eq. (8), yet it fulfills the exact linear scaling conditions (collinear or non-collinear). An extended analysis is provided in the SM.
[…]
Table 1: Spin moments (in $\mu_{b}$/atom) and, within parentheses, magnetic energies (in meV) of FM metals as calculated with DFAs and compared against experiments.
Fe Co Ni
LDA 2.23 (467) 1.57 (396) 0.56 (40.8)
PBE 2.26 (581) 1.59 (504) 0.56 (49.0)
SCAN 2.62 (1177) 1.74 (1175) 0.66 (113)
mSCAN 2.15 (461) 1.58 (447) 0.56 (47.3)
EXP 1.98-2.08Chen et al. (1995); Reck and Fry (1969); Scherz (2004) 1.52-1.62Chen et al. (1995); Reck and Fry (1969); Moon (1964); Scherz (2004) 0.52-0.55Reck and Fry (1969); Scherz (2004)
Next, we consider also materials. In Table 1 we list the spin magnetic moments (calculated analytically in the Gaussian basis set) and the magnetic energies (in meV), calculated as energy differences w.r.t. the self-consistent spin-restricted solution for the FM metals Fe, Co and Ni. All calculations employ the experimental geometries reported in Ref. Tran et al., 2020, unless explicitly stated otherwise. Comparisons are provided against the LDA (employing the parametrization of correlation energies by Vosko, Wilk and Nusair, and the usual Slater-Dirac energy functional for exchange SVWN5),Vosko et al. (1980) the GGA of PBE,Perdew et al. (1996) and the original spin-unrestricted SCAN.Sun et al. (2015) Comparison against experimental data for the spin-magnetic moments are also given.
In accord with previous reports,Fu and Singh (2018); Tran et al. (2020); Mejía-Rodríguez and Trickey (2019) we find an agreement with experiments at the level of LDA, which is not significantly changed with the GGA. SCAN displays the previously reported overmagnetization in all FM metals, both in its significant overestimation of the magnetic moments, as well as overestimation of the magnetic energies.Fu and Singh (2018); Tran et al. (2020); Mejía-Rodríguez and Trickey (2019) In the case of Fe, for example, the self-consistently predicted magnetic moment of 2.62 $\mu_{b}$ by SCAN represents an overestimation of around 30 %. Overmagnetization of around 10 % and 15 % w.r.t. experimental values are predicted by SCAN for Co and Ni respectively, while the LDA and GGA both fall within the experimental uncertainty. Magnetic energies of SCAN are consistently more than double the values predicted by the LDA and GGA for the FM metals. The problem is completely solved by switching to our gauge-invariant mSCAN approach, which provides similar magnetic moments (and magnetic energies) as the LDA and GGA on Co and Ni, while the predicted magnetic moment on Fe of 2.15 $\mu_{b}$ is the closest to the experimental range of 1.98-2.08 $\mu_{b}$.
The improvement provided by our approach is even more striking when applied to the non-magnetic metals V and Pd (see Table 2), as well as the difficult case of Cr2.Larsson et al. (2022); Zhang et al. (2020a) We perform the calculations starting from a spin-polarized guess and let the spin moments relax during the self-consistent procedure. Like in previous works,Fu and Singh (2018); Tran et al. (2020) SCAN calculations incorrectly predict a net magnetic moment of 0.33 and 0.44 $\mu_{b}$ on V and Pd respectively. In the case of Pd, the incorrect magnetic behaviour is also predicted by PBE. For Cr2, the LDA, PBE and SCAN all predict a spurious magnetization at the equilibrium geometry (ground state properties are reported in Table S1 of the SM). Remarkably, the correct non-magnetic behaviour of V, Pd and Cr2 is predicted by the proposed gauge-invariant mSCAN approach. As for itinerant two-dimensional systems, in Graphene, SCAN incorrectly predicts a magnetic insulating phase,Zhang et al. (2020a) while mSCAN correctly provides a non-magnetic, metallic state, with a Dirac point at the zone corners (Table S2).
Table 2: Same as in Table 1 but for non-magnetic systems.
V Pd Cr2
LDA 0.00 (0.00) 0.00 (0.00) 2.69 (0.30)
PBE 0.00 (0.00) 0.20 (0.70) 5.32 (0.87)
SCAN 0.57 (6.27) 0.44 (17.2) 5.62 (10.3)
mSCAN 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
EXP 0.00Weber et al. (1997) 0.00Kudasov and Korshunov (2007) 0.00Larsson et al. (2022)
00footnotetext: Values in eV.
Although the failures of SCAN on itinerant magnets are daunting, accurate magnetic moments and gaps have been predicted in the AFM Mott insulators FeO, CoO and NiO, which exhibit more localized magnetism.Zhang et al. (2020b); Tran et al. (2020) Values obtained from our mSCAN on these highly-correlated Mott insulators are provided in Table 3, including a comparison against the experiment both on magnetic moments and fundamental band gaps (in parenthesis, given in eV). The LDA and GGA incorrectly predict a metallic behaviour for FeO and CoO, while both SCAN and mSCAN provide the qualitatively correct insulating behaviour. All functionals provide similar values for the spin magnetic moments in the Mott insulators, although both moments and band gaps are larger with SCAN. We also consider a topological AFM – the Chern magnet TbMn6Sn6, in which calculation of spin moments is a reported challenge for SDFT.Annaberdiyev et al. (2023) With mSCAN, we obtain excellent agreement against the experimental values (with large improvement over SCAN and PBE) in Table S3.
Table 3: Spin moments (in $\mu_{b}$/atom) and, within parenthesis, band gaps (in eV) of AFM Mott insulators.
FeO CoO NiO
LDA 3.51 (0.000) 2.40 (0.000) 1.20 (0.452)
PBE 3.57 (0.000) 2.48 (0.000) 1.37 (0.986)
SCAN 3.67 (0.635) 2.65 (1.29) 1.62 (2.66)
mSCAN 3.57 (0.224) 2.52 (0.440) 1.44 (1.29)
EXP 2.32-4.00Roth (1958); Battle and Cheetham (1979); Fjellvåg et al. (1996) 1.75-2.98Roth (1958); Khan and Erickson (1970); Herrmann-Ronzaud et al. (1978); Jauch et al. (2001) 1.45-1.90Roth (1958); Fernandez et al. (1998); Cheetham and Hope (1983); Neubeck et al. (1999)
EXP (2.40)Bowen et al. (1975) (3.10-4.10)Gvishi and Tannhauser (1970) (3.7)Ksendzov and Drabkin (1965)
00footnotetext: We remove the orbital contributions provided in Ref. Tran et al., 2020 from the experimental total moments (see caption of Table VI of Ref. Tran et al., 2020).
Next, we investigate the performance of mSCAN on structural optimizations in all tested systems (see Table S4). A previous attempt at partially fixing the spin moments of SCAN by varying the $c_{1x}$ parameter has produced significantly worsened results for lattice constants, with errors increased by around 50%.Tran et al. (2020) Here, mean absolute errors (MAE) decrease in the order LDA, GGA, mSCAN (SCAN), with values of 0.663, 0.521, 0.295 (0.319), which is again consistent with the principle of the Jacob’s ladder of DFT.
Let us now scrutinize the dissociation of first-row diatomic molecules involving open-shell atoms.not (g) Fig. S3 reports the case of H2 compared to the accurate result of Ref. Kolos and Wolniewicz, 1968: the mSCAN and SCAN energies match up until an intermolecular separation of about 3 Å, at which point a spurious magnetization appears in the spin-unrestricted LDA, PBE and SCAN solutions. mSCAN provides a correct vanishing spin magnetization up to about 3.5 Å. In the C2 molecule LDA, PBE and SCAN provide spurious magnetization even at the equilibrium geometry, which is resolved by mSCAN, see Fig. S4 and Table S5. mSCAN also provides an improvement over SCAN, with respect to the reference total energies of Ref. Booth et al., 2011 on C2, both at the equilibrium geometry and at large distances. Fig. S5 reports the dissociation of the N2 molecule. mSCAN provides a significant improvement over LDA, PBE and SCAN at distances larger than about 1.5 Å, where strong correlation becomes important (as also corroborated by comparison of multi- and single-reference coupled-cluster treatments).Li and Paldus (2008)
A word of caution and of encouragement are in order: The novel functional behaves differently from SCAN not only for magnetized states but also for open-shell states, therefore atomization energies should be scrutinized with care. We have also pointed out that band gaps may be worsened. Therefore, superseding the SCAN functional in all respects will require further efforts, which, we believe, can exploit the recently discovered extra-degrees of freedom that are allowed by the MGGAs. Lebeda et al. (2024)
Figure 1: Magnetization in $\mu_{b}$ of the Cr3 molecule with the nc SCAN functional.
Noncollinear extension.  Fully self-consistent SCDFT MGGA calculations for non-collinear magnetic states are reported in a parallel work.Huebsch et al. (2025) Ref. Huebsch et al., 2025 also calls for a gauge-invariant nc-extension of SCAN.
Thus, we finally solve Eq. (10) using the nc forms in Eq.s (13a) and (13b). Technical details are provided in the SM, where we also demonstrate that the functional provides a non-zero xc magnetic torque, which vanishes over the full space. Last but not least, spin-orbit coupling can be accounted for self-consistently. So, our proposal fulfills the most pressing desiderata for a nc approach.
We applied it to the Cr3 (bond length of 3.7 Bohr), which has been an important model system in previous studies.Scalmani and Frisch (2012); Desmarais et al. (2024a); Huebsch et al. (2025); Tancogne-Dejean et al. (2023); Pu et al. (2023) The magnetization reported in Fig. 1 is indicative of a spin-frustrated system. We observe a spin-moment of 1.38 $\mu_{b}$, which is significantly lower than the SCDFT-MGGA value of 3.15 $\mu_{b}$ reported in Ref. Huebsch et al., 2025, but closer to the LDA value of 1.68 $\mu_{b}$.
Conclusions.   While it remains a challenge to assess whether a semi-local density functional may be derived to embrace all the relevant non-magnetic and magnetic (both collinear and non-collinearly polarized) exact constraints, the core findings of this work solves most urgent failures for magnetism and resumes progress towards improved density functional approximations.
Acknowledgements.
This research has received funding from the Project CH4.0 under the MUR program “Dipartimenti di Eccellenza 2023-2027” (CUP: D13C22003520001). GV was supported by the Ministry of Education, Singapore, under its Research Centre of Excellence award to the Institute for Functional Intelligent Materials (I-FIM, project No. EDUNC-33-18-279-V12).